TWO-TOP TORSIONAL ASSIGNMENTS IN THE A^{1}A_{u}(S_{1}) − X^{1}A_{g}(S_{0}) TRANSITION OF BIACETYL CH_{3}C(=O)C(=O)CH_{3}

C.-L. HUANG, Department of Applied Chemistry, National Chiayi University, Chiayi, Taiwan; C.-L. LIU, C.-K. NI, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 106, P.O. Box 23-166, Taiwan; \underline{J. T. HOUGEN}, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The laser-induced fluorescence excitation spectrum of the biacetyl A^{1}A_{u}(S_{1}) − X^{1}A_{g}(S_{0}) transition in the region from 22 182 to 28 000 cm^{−1}shows a complicated absorption line spectrum, which is believed to arise from a long progression in the torsional vibrations of the two equivalent methyl tops in this molecule. In this talk we discuss three topics: (i) a numerical calculation of these energy levels using a kinetic and potential energy formalism and preliminary constants from the literature,^{a} (ii) a qualitative description of the calculated energy level pattern using local mode ideas applied to the two equivalent methyl rotors as well as G_{36} permutation-inversion group symmetry species, and (iii) a comparison of the calculated energy levels with the low-resolution spectrum taken in the region from 0 to 500 cm^{−1}above the A-X band origin at IAMS in Taipei. The levels with one quantum of torsional excitation are best described by a normal mode formulation, but levels with more than one quantum of torsional excitation are better described by a local mode formulation. Good agreement with the low resolution spectrum is obtained, but full confirmation of the interpretation will require high resolution studies of some additional bands.

aM. L. Senent, D. C. Moule, Y. G. Smeyers, A. Toro-Labbé, and F. J. Peqalver, J. Mol. Spectrosc. 164, 66-78 (1994).